# Large Language Models for Scholarly Ontology Generation: An Extensive Analysis in the Engineering Field

### v1

@@ 3.2 Gold Standard
To evaluate the performance of LLMs on this task, we created a gold standard by selecting a sample of 1,000 semantic relationships from the IEEE Thesaurus. We have named this dataset IEEE-Rel-1K (IEEE Relations). The IEEE thesaurus contains around 11,730 engineering, technical, and scientific terms, including IEEE-specific society terms. It is compiled from IEEE transactions, journal articles, conference papers, and standards, and reflects the vocabulary used within the engineering and scientific fields of IEEE. Specifically, we utilised the IEEE Thesaurus v1.02444A copy of version 1.02 of the IEEE Thesaurus is available at https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf/blob/main/source/ieee-thesaurus_2023.pdf., which is dated back to July 2023, and is available as a PDF following the ANSI/NISO Z39.4-2021 standard [31].
To create IEEE-Rel-1K, we developed a Python script555Script for extracting semantic relationships in IEEE Thesaurus - https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf. for extracting the hierarchical structure and relationships between terms from the IEEE Thesaurus PDF file. Next, we randomly selected 250 relationships for each of the categories: broader, narrower, and same-as. For same-as, we used a combination of ‘use preferred term’ and ‘used for’ relationships as defined in the ANSI/ISO standard. To complete the dataset, we randomly generated 250 pairs of topics, ensuring they were not semantically related within the IEEE Thesaurus, and labelled these as other.
@@ 3.3.1 Open Models
Mistral-7B (shortened as mistral) is trained on Open-Web data, has 7 billion parameters, and utilises grouped-query attention and sliding window attention with a context window of 32,000 tokens.
Mixtral-8×7B (shortened as mixtral) is a sparse mixture-of-experts network incorporating 46.7 billion parameters. Only 12.9 billion parameters are used per token, whereas the remaining parameters are part of the “expert” layers within its sparse mixture-of-experts architecture. It has a context window of 32,000 tokens.
Llama-2-70B (shortened as llama-2) has 70 billion parameters and a context window of 4,096 tokens. It is trained on a corpus that includes 2 trillion tokens of data from publicly available sources.
@@ 3.3.2 Quantised Models
Dolphin-2.1-Mistral-7B (shortened as dolphin-mistral666Dolphin-2.1-Mistral-7B - https://huggingface.co/TheBloke/dolphin-2.1-mistral-7B-GGUF) is a model with 7 billion parameters and a token context capacity of 4,096. It is based on Mistral-7B and fine-tuned with the Dolphin777Dolphin Dataset - https://huggingface.co/datasets/cognitivecomputations/dolphin dataset, which is an open-source implementation of Microsoft’s Orca [32], with an addition of Airoboros888Airoboros - https://huggingface.co/datasets/jondurbin/airoboros-2.1 dataset.
@@
Dolphin-2.6-Mistral-7B-dpo-laser (shortened as dolphin-mistral-dpo999Dolphin-2.6-Mistral-7B-dpo-laser - https://huggingface.co/TheBloke/dolphin-2.6-mistral-7B-dpo-laser-GGUF) is based on Mistal-7B with 7 billion parameters and fine-tuned on top of Dolphin DPO using Layer Selective Rank Reduction (LASER) [33]. It offers a context window of 4,096 tokens.
@@
Dolphin2.1-OpenOrca-7B (shortened as dolphin-openorca101010 Dolphin2.1-OpenOrca-7B - https://huggingface.co/TheBloke/Dolphin2.1-OpenOrca-7B-GGUF) is a model that blends Dolphin-2.1-Mistral-7B and Mistral-7B-OpenOrca models. These models were merged using the “ties merge” [34] technique, keeping the same number of training parameters and token context window size, respectively 7 billion and 4,096.
@@
OpenChat-3.5-1210 (shortened as openchat111111OpenChat-3.5-1210 - https://huggingface.co/TheBloke/openchat-3.5-1210-GGUF) is a model fine-tuned on top of Mistral-7B-v0.1. It possesses similar properties and consists of 7 billion parameters with a context window size of 8,192 tokens.
@@
OpenChat-3.5-0106-Gemma (shortened as Openchat-gemma121212OpenChat-3.5-0106-Gemma - https://huggingface.co/gguf/openchat-3.5-0106-gemma-GGUF) is a model trained on openchat-3.5-0106131313openchat-3.5-0106 - https://huggingface.co/openchat/openchat-3.5-0106 data using Conditioned Reinforcement Learning Fine-Tuning (C-RLFT) framework [35, 36]. This model shares the same properties of the openchat-3.5-0106 model, which was fine-tuned141414https://huggingface.co/openchat/openchat-3.5-1210/discussions/4#658288f1168803bdee13d6b3 on top of Mistral-7B. It consists of 7 billion parameters with a context window of 8,192 tokens.
@@
SOLAR-10.7B-Instruct-v1.0 (shortened as solar151515SOLAR-10.7B-Instruct-v1.0 - https://huggingface.co/TheBloke/SOLAR-10.7B-Instruct-v1.0-GGUF) a 10.7 billion parameters model with a Depth Up-Scaling (DUS) architecture that includes architectural modifications and continued pretraining [37]. It is trained on the OpenOrca161616OpenOrca dataset - https://huggingface.co/datasets/Open-Orca/OpenOrca dataset.
@@
Mistral-7B-OpenOrca (shortened as mistral-openorca171717Mistral-7B-OpenOrca - https://huggingface.co/Open-Orca/Mistral-7B-OpenOrca) is based on Mistral-7B-Instruct-v0.1 and has been fine-tuned using the OpenOrca dataset [32]. This dataset is a comprehensive collection of augmented Fine-tuned LAnguage Net (FLAN) data aligned with the distributions outlined in the Orca paper [38, 39]. The model has 7 billion parameters and a context window of 4,096.
@@
Eurus-7b-sft (shortened as eurus181818Eurus-7b-sft - https://huggingface.co/mradermacher/Eurus-7b-sft-GGUF) is fine-tuned on top of Mistral-7B-Instruct-v0.2 using all correct actions in UltraInteract, mixing a small proportion of UltraChat, ShareGPT, and OpenOrca examples, and posses 7 billion parameters with a context window size of 4,096 [40].
@@
Llama-3-8B-Instruct (shortened as llama-3191919Llama-3-8B-Instruct - https://huggingface.co/lmstudio-community/Meta-Llama-3-8B-Instruct-GGUF) has 8 billion parameters and a context length of almost 8,192 tokens [41]. It utilises supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
@@
Orca-2-13B (shortened as orca-2202020Orca-2-13B - https://huggingface.co/TheBloke/Orca-2-13B-GGUF) is a model with 13 billion parameters and a 2,048-token context window. It was trained on a mix of publicly available and curated datasets to enhance its reasoning and comprehension abilities. The fine-tuning process was based on datasets like Super-NaturalInstructions and FLAN to ensure strong performance across NLP tasks [42].
@@ 3.3.3 Proprietary Models
GPT-3.5-Turbo-Instruct (shortened as gpt-3.5) is trained using RLHF, following similar methods as InstructGPT, but with some differences in the data collection setup. It has a context window size of 4,096 tokens and was trained with data up to September 2021.
GPT-4-Turbo (shortened as gpt-4) has been pre-trained to predict the next token in a document. It was trained using publicly available data and data licensed from third-party providers up to December 2023. The model was then fine-tuned using RLHF. It has a context window of 128,000 tokens.
Claude 3 Haiku (shortened as haiku) is a model with a context size of 200,000 tokens, offering a maximum output of 4,096 tokens with training data up to August 2023.
@@ 4 Experiments
We evaluated the 17 models described in Section 3.3 against IEEE-Rel-1K, the gold standard defined in Section 3.2, employing four different experiments that varied based on the prompting strategy utilised. Figure 1 illustrates the flowcharts of these experiments.
The experiments were categorised into two primary groups: standard prompting and chain-of-thought (CoT) prompting [44]. For each category, we conducted two distinct experiments. The first approach, labelled one-way, involves the model inferring the relationship between and directly. The second approach, referred to as two-way, builds upon the first by having the model initially infer the relationship between and , then infer the relationship between and , and finally combine the results using heuristic rules.
In all cases, the performance of the model was assessed using standard precision, recall and F1.
To summarise, the four experiments can be labelled as follows:
In the following, we will describe them in detail.
@@ 4.2 Chain-of-thought prompting
CoT prompting experiments operated similarly to the previous solution but involved a two-phase interaction with the LLM, as illustrated in Fig. 1(b). In the first phase, the model is asked to provide a definition of both topics, formulate a sentence that incorporates both of them, and finally discuss their potential semantic relationships. The resulting response is then fed back into the model in the second phase, with the sole task of determining the type of semantic relationship. The two templates used to produce the prompts for the chain-of-thought experiments are available in B.
@@
The experiments for both standard prompting and CoT prompting are conducted in two different configurations. The one-way strategy simply asks for the relationship between and . The two-way strategy involves asking the LLM to identify the relationship between topics and , as well as the relationship between topics and , when the order of the topics reversed. This straightforward approach helps to detect and potentially correct inconsistencies in the model’s responses. This is feasible because the broader and narrower are inverse relationships, while same-as and other are symmetric. For example, if the model determines that is broader than , but also that is same-as , we can infer that at least one of these answers is incorrect.
To manage these potential inconsistencies, we designed a set of empirical rules (cyan boxes as referees in both Fig. 1(a) and 1(b)). These rules were designed to prioritise the development of the hierarchical structure.
@@ 4.4 Experimental setup
To interact with the various LLMs we employed two services, e.g., Amazon Bedrock212121Amazon Bedrock - https://aws.amazon.com/bedrock, and OpenAI API222222OpenAI API - https://platform.openai.com/docs/overview; as well as the KoboldAI232323Kobold API - https://github.com/KoboldAI/KoboldAI-Client library for running LLMs locally.
@@ 5 Results and Discussion
This section presents the results of the complete set of experiments described in Section 4. First we will discuss the standard prompting experiments (see Fig. 1(a) for flowchart) and then we will report the results of the chain-of-thought prompting experiments (see Fig. 1(b)). Further, we will compare the results across the various strategies.
@@ 5.1 Standard prompting, one-way strategy
Table 1 reports the result of the experiments with standard prompting when applying one-way strategy. Among the full open models (first three rows), llama-2 and mixtral demonstrate good performance in terms of average F1 with 0.669 and 0.779, respectively. In particular, mixtral is highly precise with the broader (0.819) and same-as (1), whereas llama-2 obtains high values of precision with narrower (1) and other (0.877). Complementarily, mixtral reaches high values of recall for narrower (0.888) and other (0.940), whereas llama-2 for broader (0.872) and same-as (0.992). In this case, mistral consistently performs low, except for the recall of the relationship other (0.992).
Among quantised models, dolphin-openorca and openchat stand out with the highest average F1-score of 0.724 and 0.705, respectively. While both consistently excel in precision, they face some challenges in recall for narrower and same-as relations. Other quantised models, although falling below 0.7 in average F1-score, still demonstrate remarkable performance in specific relationships. For instance, mistral-openorca excels in precision for the same-as (1) and other (0.992) relations and recall for narrower (0.996), despite lower performance in all other metrics.
When evaluating the proprietary models (as shown in the last four rows of Table 1), sonnet stands out, achieving an outstanding average F1-score of 0.967, with a well-balanced precision and recall across all relationship types. gpt-4 also demonstrates strong performance (0.952 F1). These two models exhibit high precision for same-as and other relations, and high recall for broader relations. In contrast, the two other models do not perform as well. Haiku, with an F1-score of 0.712, is surpassed by both mixtral and dolphin-openorca. Surprisingly, gpt-3.5 also shows lower-than-expected results.
Overall, we observe a significant performance gap between full open models and proprietary models. For example, sonnet outperforms mixtral, achieving an average F1-score that is 0.188 points higher.

### v2

@@ 3.2 Gold Standard
To evaluate the performance of LLMs on this task, we created a gold standard by selecting a sample of 1,000 semantic relationships from the IEEE Thesaurus. We have named this dataset IEEE-Rel-1K (IEEE Relations). The IEEE thesaurus contains around 11,730 engineering, technical, and scientific terms, including IEEE-specific society terms. It is compiled from IEEE transactions, journal articles, conference papers, and standards, and reflects the vocabulary used within the engineering and scientific fields of IEEE. Specifically, we utilised the IEEE Thesaurus v1.02555A copy of version 1.02 of the IEEE Thesaurus is available at https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf/blob/main/source/ieee-thesaurus_2023.pdf., which is dated back to July 2023, and is available as a PDF following the ANSI/NISO Z39.4-2021 standard [56].
To create IEEE-Rel-1K, we developed a Python script666Script for extracting semantic relationships in IEEE Thesaurus - https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf. for extracting the hierarchical structure and relationships between terms from the IEEE Thesaurus PDF file. Next, we randomly selected 250 relationships for broader and narrower. For same-as, we used a combination of ‘use preferred term’ and ‘used for’ relationships as defined in the ANSI/NISO Z39.19-2005 standard [22]. While the IEEE Thesaurus offers numerous related topics, these do not consistently meet the strict definition of synonymy [22]. For instance, ‘4G mobile communication’ is related to ‘5G mobile communication’ but while they belong to the same category of mobile networks, they represent entirely different technologies. Therefore, a team of three experts manually validated and selected only synonyms, lexical variants, and near-synonyms. This manual validation was time-consuming, limiting our dataset to 250 samples per category. To complete the dataset, we randomly generated 250 pairs of topics, ensuring they were not semantically related within the IEEE Thesaurus, and labelled these as other.
@@ 3.3.1 Open Models
Mistral-7B (shortened as mistral) is trained on Open-Web data, has 7 billion parameters, and utilises grouped-query attention and sliding window attention with a context window of 32,000 tokens [57].
Mixtral-8×7B (shortened as mixtral) is a sparse mixture-of-experts network incorporating 46.7 billion parameters. Only 12.9 billion parameters are used per token, whereas the remaining parameters are part of the “expert” layers within its sparse mixture-of-experts architecture. It has a context window of 32,000 tokens [58].
Llama-2-70B (shortened as llama-2) has 70 billion parameters and a context window of 4,096 tokens. It is trained on a corpus that includes 2 trillion tokens of data from publicly available sources [59].
@@ 3.3.2 Quantised Models
Dolphin-2.1-Mistral-7B (shortened as dolphin-mistral777Dolphin-2.1-Mistral-7B - https://huggingface.co/TheBloke/dolphin-2.1-mistral-7B-GGUF) is a model with 7 billion parameters and a token context capacity of 4,096. It is based on Mistral-7B and fine-tuned with the Dolphin888Dolphin Dataset - https://huggingface.co/datasets/cognitivecomputations/dolphin dataset, which is an open-source implementation of Microsoft’s Orca [60], with an addition of Airoboros999Airoboros - https://huggingface.co/datasets/jondurbin/airoboros-2.1 dataset.
@@
Dolphin-2.6-Mistral-7B-dpo-laser (shortened as dolphin-mistral-dpo101010Dolphin-2.6-Mistral-7B-dpo-laser - https://huggingface.co/TheBloke/dolphin-2.6-mistral-7B-dpo-laser-GGUF) is based on Mistal-7B with 7 billion parameters and fine-tuned on top of Dolphin DPO using Layer Selective Rank Reduction (LASER) [61]. It offers a context window of 4,096 tokens.
@@
Dolphin2.1-OpenOrca-7B (shortened as dolphin-openorca111111 Dolphin2.1-OpenOrca-7B - https://huggingface.co/TheBloke/Dolphin2.1-OpenOrca-7B-GGUF) is a model that blends Dolphin-2.1-Mistral-7B and Mistral-7B-OpenOrca models. These models were merged using the “ties merge” [62] technique, keeping the same number of training parameters and token context window size, respectively 7 billion and 4,096.
@@
OpenChat-3.5-1210 (shortened as openchat121212OpenChat-3.5-1210 - https://huggingface.co/TheBloke/openchat-3.5-1210-GGUF) is a model fine-tuned on top of Mistral-7B-v0.1. It possesses similar properties and consists of 7 billion parameters with a context window size of 8,192 tokens [63].
@@
OpenChat-3.5-0106-Gemma (shortened as Openchat-gemma131313OpenChat-3.5-0106-Gemma - https://huggingface.co/gguf/openchat-3.5-0106-gemma-GGUF) is a model trained on openchat-3.5-0106141414openchat-3.5-0106 - https://huggingface.co/openchat/openchat-3.5-0106 data using Conditioned Reinforcement Learning Fine-Tuning (C-RLFT) framework [64, 63]. This model shares the same properties of the openchat-3.5-0106 model, which was fine-tuned151515https://huggingface.co/openchat/openchat-3.5-1210/discussions/4#658288f1168803bdee13d6b3 on top of Mistral-7B. It consists of 7 billion parameters with a context window of 8,192 tokens.
@@
SOLAR-10.7B-Instruct-v1.0 (shortened as solar161616SOLAR-10.7B-Instruct-v1.0 - https://huggingface.co/TheBloke/SOLAR-10.7B-Instruct-v1.0-GGUF) a 10.7 billion parameters model with a Depth Up-Scaling (DUS) architecture that includes architectural modifications and continued pretraining [65]. It is trained on the OpenOrca171717OpenOrca dataset - https://huggingface.co/datasets/Open-Orca/OpenOrca dataset.
@@
Mistral-7B-OpenOrca (shortened as mistral-openorca181818Mistral-7B-OpenOrca - https://huggingface.co/Open-Orca/Mistral-7B-OpenOrca) is based on Mistral-7B-Instruct-v0.1 and has been fine-tuned using the OpenOrca dataset [60]. This dataset is a comprehensive collection of augmented Fine-tuned LAnguage Net (FLAN) data aligned with the distributions outlined in the Orca paper [66, 67]. The model has 7 billion parameters and a context window of 4,096.
@@
Eurus-7b-sft (shortened as eurus191919Eurus-7b-sft - https://huggingface.co/mradermacher/Eurus-7b-sft-GGUF) is fine-tuned on top of Mistral-7B-Instruct-v0.2 using all correct actions in UltraInteract, mixing a small proportion of UltraChat, ShareGPT, and OpenOrca examples, and posses 7 billion parameters with a context window size of 4,096 [68].
@@
Llama-3-8B-Instruct (shortened as llama-3202020Llama-3 - https://huggingface.co/lmstudio-community/Meta-Llama-3-8B-Instruct-GGUF) has 8 billion parameters and a context length of almost 8,192 tokens [69]. It utilises supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
@@
Orca-2-13B (shortened as orca-2212121Orca-2-13B - https://huggingface.co/TheBloke/Orca-2-13B-GGUF) is a model with 13 billion parameters and a 2,048-token context window. It was trained on a mix of publicly available and curated datasets to enhance its reasoning and comprehension abilities. The fine-tuning process was based on datasets like Super-NaturalInstructions and FLAN to ensure strong performance across NLP tasks [70].
@@ 3.3.3 Proprietary Models
GPT-3.5-Turbo-Instruct (shortened as gpt-3.5) is trained using RLHF, following similar methods as InstructGPT, but with some differences in the data collection setup. It has a context window size of 4,096 tokens and was trained with data up to September 2021 [71].
GPT-4-Turbo (shortened as gpt-4) has been pre-trained to predict the next token in a document. It was trained using publicly available data and data licensed from third-party providers up to December 2023. The model was then fine-tuned using RLHF. It has a context window of 128,000 tokens [72].
Claude 3 Haiku (shortened as haiku) is a model with a context size of 200,000 tokens, offering a maximum output of 4,096 tokens with training data up to August 2023 [73].
@@ 4 Experiments
We evaluated the 17 models described in Section 3.3 against IEEE-Rel-1K, the gold standard defined in Section 3.2, employing four different experiments that varied based on the prompting strategy utilised.
We investigated two primary prompting paradigms: a standard prompting approach based on a simple prompt that directly asks the model to perform the classification and Chain-of-Thought (CoT) prompting [74], which encourages the model to reason through a predefined sequence of steps before providing an answer. For each of these two paradigms, we employed two additional strategies, resulting in four distinct experimental settings. The first, labelled one-way, involves the model inferring the relationship between and directly. The second, referred to as two-way, builds upon the first by having the model initially infer the relationship between and , then infer the relationship between and , and finally combine the results using heuristic rules.
Figure 1 illustrates the flowcharts of these experiments.
Based on these strategies, we can characterise our four experiments as follows:
In all experiments, the performance of the model was assessed using standard precision, recall and F1.
The standard prompting, one-way strategy serves as a baseline for classification tasks. We introduced CoT prompting due to its demonstrated effectiveness in complex reasoning tasks [74, 75]. Indeed, as discussed in Section 5, CoT prompting led to improved performance for most models. We investigated the two-way validation strategy to identify and rectify potential inconsistencies. It draws inspiration from prior research showing that approaches to improve logical consistency can enhance the overall performance of LLMs [76]. This approach leverages the inverse relationship between broader and narrower, as well as the symmetric nature of the same-as and other relationships [22]. Conflicting answers suggest inconsistencies that may be addressed using heuristic methods.
In the following subsections, we provide a detailed description of each strategy, followed by a discussion of the experimental setup and implementation details.
@@ 4.2 Chain-of-thought prompting
CoT prompting experiments operated similarly to the previous solution but involved a two-phase interaction with the LLM, as illustrated in Fig. 1(b). In the first phase, the model is asked to provide a definition of both topics, formulate a sentence that incorporates both of them, and finally discuss their potential semantic relationships. The resulting response is then fed back into the model in the second phase, with the sole task of determining the type of semantic relationship. The two templates used to produce the prompts for the CoT experiments are available in B.
@@
The experiments for both standard prompting and CoT prompting are conducted in two different configurations. The one-way strategy simply asks for the relationship between and . The two-way strategy involves asking the LLM to identify the relationship between topics and , as well as the relationship between topics and , where the order of the topics is reversed. This approach can naturally produce inconsistent responses, which may indicate potential issues.
In order to detect and mitigate logical inconsistencies, we designed a set of empirical rules (cyan boxes as referees in both Fig. 1(a) and 1(b)). The rules were designed to prioritise the development of a hierarchical taxonomy.
@@ 4.4 Experimental setup
To interact with the various LLMs we employed two services, e.g., Amazon Bedrock222222Amazon Bedrock - https://aws.amazon.com/bedrock, and OpenAI API232323OpenAI API - https://platform.openai.com/docs/overview; as well as the KoboldAI242424Kobold API - https://github.com/KoboldAI/KoboldAI-Client library for running LLMs locally.
@@ 5 Results and Discussion
This section presents the results of the complete set of experiments described in Section 4. First we will discuss the standard prompting experiments (see Fig. 1(a) for flowchart) and then we will report the results of the CoT prompting experiments (see Fig. 1(b)). Further, we will compare the results across the various strategies.
@@ 5.1 Standard prompting, one-way strategy
Table 1 reports the results of the experiments with standard prompting when applying one-way strategy. Among the full open models (first three rows), llama-2 and mixtral demonstrate good performance in terms of average F1 with 0.669 and 0.779, respectively. In particular, mixtral shows a high value of precision with the broader (0.819) and same-as (1), whereas llama-2 obtains high values of precision with narrower (1) and other (0.877). Complementarily, mixtral reaches high values of recall for narrower (0.888) and other (0.940), whereas llama-2 for broader (0.872) and same-as (0.992). In this case, mistral consistently performs low, except for the recall of the relationship other (0.992). The observed performance differences may be partially attributed to variations in the pre-training datasets. Indeed, mixtral [57] was trained on a more diverse corpus that included scientific and technical documents [58]. This broader exposure may have enhanced its ability to model the relationships between scientific domains, resulting in superior performance in identifying hierarchical structures. In contrast, llama-2 [59] was pre-trained primarily on web data.
Among quantised models, dolphin-openorca and openchat stand out with the highest average F1-score of 0.724 and 0.705, respectively. While both consistently excel in precision, they face some challenges in recall for narrower and same-as relations. Other quantised models, although falling below 0.7 in average F1-score, still demonstrate remarkable performance in specific relationships. For instance, mistral-openorca excels in precision for the same-as (1) and other (0.992) relations and recall for narrower (0.996), despite lower performance in all other metrics. The varied performance among models with the same base architecture but different fine-tuning approaches highlights the critical role of instruction-tuning datasets [77]. dolphin-openorca and openchat were both fine-tuned on datasets that emphasise following detailed instructions [60, 63], which may make them more suitable for our specific task.
When evaluating the proprietary models (as shown in the last four rows of Table 1), sonnet stands out, achieving an outstanding average F1-score of 0.967, with a well-balanced precision and recall across all relationship types. gpt-4 also demonstrates strong performance (0.948 F1). These two models exhibit high precision for same-as and other relations, and high recall for broader relations. In contrast, the two other models do not perform as well. Haiku, with an F1-score of 0.712, is surpassed by both mixtral and dolphin-openorca. Surprisingly, gpt-3.5 also shows lower-than-expected results. The exceptional performance of sonnet and gpt-4 is most likely attributable to both their size and their pre-training on datasets that include extensive scientific and technical corpora [73, 72]. While the specific composition of their training data remains proprietary, these models have demonstrated superior capabilities on several tasks that require a nuanced understanding of scientific domains [78]. Naturally, both model size and recency play a significant role, as evidenced by the superior performance of the most recent and largest proprietary models (gpt-4, sonnet) compared to their earlier and smaller counterparts (gpt-3.5, haiku).
Overall, we observe a clear performance gap between full open models and proprietary models. For example, sonnet achieves an average F1-Score of 0.967, substantially outperforming mixtral which has an average F1-Score of 0.779.
While none of the models explicitly claim to have been trained on the IEEE digital library specifically, the superior performance of proprietary models suggests they may have encountered scientific publications during pre-training [79]. Kandpal et al. [80] demonstrated that LLMs can effectively memorise portions of their training data, particularly for well-structured technical content.